Accelerator Update for March 12, 2008 - March 14, 2008
Wednesday March 12
The day shift began with Operations monitoring the Tevatron (TeV) store 5977 (92.56E30) and the Recycler stash (188.9E10), with the Antiproton Source stacking (39.1E10), with the Main Injector (MI), Linac, and Booster operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with NuMI and the SciBooNE / MiniBooNE (BNB) experiments taking beam.
At 8:20 AM, LRF3 tripped off. It reset.
RF experts reported at 10:12 AM that they were investigating intensity problems at MI-60.
Operations prepared NuMI for an access at 10:31 AM.
At 12:33 PM, Operations reported that the Pbar sweeping magnets were tripping off. Controls for these magnets came through a Pbar PC (PBKICK) that was recently rebooted. An expert discovered that bad settings were loaded during the reboot.
At 12:56 PM, KRF3 tripped off. It reset.
RF experts began work on MIRF11 at 1:37 PM.
At 3:04 PM, KRF2 tripped off. It reset.
Operations reported at 3:58 PM that the Recycler had stashed antiprotons five times during the day shift.
At 4:19 PM, the Linac Klystron Vernier tripped off on a transformer LCW flow fault. It reset.
Operations terminated store 5977 at 5:06 PM.
Operations put the TeV into shot setup at 5:17 PM.
A MI expert reported at 6:10 PM that their collimation orbit study was successful. They will want to run this study again. He will leave the collimators set at their standard non-collimating one-inch horizontal positions and the 0.2-inch vertical positions.
Operations established store 5978 at 7:01 PM with an initial luminosity of 259.0E30.
At 8:12 PM, Operations reported that a TeV RF station (TRF8) lost its gap envelope and tripped off. This caused the loss of about 1E30 of luminosity. Operators turned off the station’s heaters.
NuMI came out of access at 8:14 PM.
At 8:26 PM, KRF4 tripped off. The station reset, but would not phase in. Operations contacted an expert who instructed operators on how to get the station up and operating, which occurred by 9:02 PM.
Linac resumed sending beam at 9:03 PM. NuMI resumed taking beam at this time.
Operations reported at 11:59 PM that the Recycler had stashed four times during the evening shift.
Thursday March 13
The midnight shift began with Operations monitoring the stash (103.5E10), with Pbar stacking (35.1E10), and with NuMI and BNB taking beam.
Operations reported at 12:27 AM that TeV sector A4 reciprocating cold compressor was inspected and found in good condition. It was restarted. Operations will run the compressor for about an hour before Operations puts the TeV into shot setup.
Operations put the TeV through a dry squeeze at 3:44 AM.
Operations put the TeV into shot setup at 5:15 AM.
At 5:39 AM, MIRF5 tripped off due to a bias power supply trip. It would not reset remotely. This trip held off the NuMI beam permit. Operators had the station back online within six minutes.
At 6:25 AM, KRF3 tripped off. It reset.
At 7:02 AM, Operations reported that the third transfer of antiprotons to the TeV were lost. About 75% of the antiprotons were lost over three transfers. Operators reported that a Recycler power supply (R:H330) was not ramping.
At 7:51 AM, a Recycler expert had tried to rig R:H330 to ramp properly, but it didn’t work. Experts conferred.
Operations reported at 8 AM that the Recycler had stashed two times during the midnight shift.
A Recycler expert reported at 8:31 AM that R:H330 was ramping. Operations transferred the remaining antiprotons.
At 8:50 AM, KRF2 tripped off. It reset.
At 9:05 AM, the TeV sector A17 separator sparked. It reset.
Operations reported at 9:10 AM that a Recycler expert continued to work on R:H330 during the transfer. Operators estimated that only 5E10 of antiprotons were lost.
Water technicians reported at 9:15 AM that they were making up water for the MI-60 LCW system.
Operations established store 5980 at 9:25 AM with an initial luminosity of 58.7E30.
RF experts began work on MIRF5 at 10:09 AM.
At 11:39 AM, experts reported finding an unusual ripple on the supply output of B0Q5. They think an SCR wasn’t firing on pulse #1. They will investigate further in between stores.
At 1:08 PM, the TeV sector F4 air compressor tripped off. It reset.
Operations reported at 2:18 PM that they had turned on three vacuum pumps in MI (IP817, IP5282, and IP5321).
At 2:36 PM, the NuMI beam permit tripped off due to wandering current on a power supply (E:V108). Everything reset.
Operations reported at 3:57 PM that the Recycler had stashed five times during the day shift.
At 4:04 PM, KRF2 tripped off. It reset.
At 4:12 PM, a Booster vacuum pump (IP62) tripped off. It reset.
Operations reported at 4:24 PM that TeV experts had been conducting end-of-store orbit smoothing studies.
Operations terminated store 5980 at 4:27 PM.
TeV power supply experts began work on B0Q5 at 5:09 PM.
At 6:27 PM, the NuMI beam permit tripped off due to beam in the kicker gap. It reset.
Cryo system technicians reported at 6:35 PM that they had completed their work on the TeV sector A4 reciprocating cold compressor and moved on to the C4 compressor.
A TeV expert requested at 7:41 PM that Operations rebooted the collimator FE at TeV sector D17.
A MI collimator expert reported at 7:48 PM that the collimation work today resulted in a new configuration that improved stashing and transfers to the TeV.
Power supply experts reported at 8:12 PM that they had completed their work on B0Q5.
Operations put the TeV through a dry squeeze at 8:53 PM.
Operations put the TeV through a wet squeeze at 9:16 PM.
At 9:43 PM, LRF2 tripped off. It reset.
At 10:54 PM, Operations reported that a number of TeV LCW alarms posted. Operators investigated and discovered that the TeV sector F4 LCW pump #2 had tripped off. It reset.
Operations put the TeV into shot setup at 11:12 PM.
Operations reported at 11:54 PM that the Recycler had stashed three times during the evening shift.
Friday March 14
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (305.8E10), with Pbar stacking (37.8E10), and with NuMI and BNB taking beam.
At 12:35 AM, KRF4 tripped off. It reset.
At 1:27 AM, KRF6 tripped off. It reset.
Operations established store 5984 at 1:31 AM with an initial luminosity of 235.9E30.
At 6:10 AM, LRF3 tripped off. It reset.
Operations put Pbar into standby at 7:46 AM.
Operations prepared MI-8 and MI-12 for access at 7:47 AM.
Operations prepared Booster for access at 7:48 AM.
Operations loaded an energy conservation time line at 7:53 AM.
Operations reported at 8 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today are to look for a possible LCW leak in Booster, MI-8 & 12 this morning and then resume normal stacking, stashing, and storing for today and the weekend.
Comments and Suggestions
|last modified 03/14/2008 email Fermilab|
|Security, Privacy, Legal|